Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by arorag » Sun Jun 29, 2008 5:36 pm
From a group of 3 boys and 3 girls, 4 children are to be randomly selected. What is the probability that equal numbers of boys and girls will be selected?
ANS[spoiler] 4/9[/spoiler]
Join the discussion
Source: — Problem Solving |

by punit.kaur.mba » Tue Jul 01, 2008 8:10 am
According to me the answer should be 9/15 = 3/5

4/9 looks wrong to me. Reason being, he is simple taking 2/3 * 2/3 = 4/9

Doesnt make sense...



Here is how I get 3/5 -
Denominator - Total no of ways to get 4 children = 6C4
Numerator - Total no of ways to select 2 girls of 3 * Total no of ways to select 2 boys of 3
=(3C2 * 3C2)/6C4
=9/15 = 3/5

arorag, could you please check again. I doubt the OA. Whats the source of the quest..

Anybody has any other explainations??
Join the discussion

by Ian Stewart » Tue Jul 01, 2008 10:51 am
Yes, the OA (4/9) can't be right here. The answer is 3/5; I used the same approach as in the above post.
Join the discussion

by beeparoo » Wed Jul 02, 2008 5:28 pm
Ian Stewart wrote:Yes, the OA (4/9) can't be right here. The answer is 3/5; I used the same approach as in the above post.
Thanks for saying something, punit, Ian! I was concerned too when I saw the answer. Phew.
Join the discussion

by somail » Thu Jul 03, 2008 10:57 am
Hi,

For some reason I am having a hard time understanding the probablity questions that are usually posted. Can one of you solve this in a little more detail (dumbed down).

Thanks for any help
Join the discussion

by pranavc » Thu Jul 03, 2008 6:38 pm
I'll give this a shot. You have three guys and three girls to select from. You have 4 spots. The order in these 4 spots is insignificant. So you have to use the formula of combinations to determine the total number of selection possibilities. This boils down to 6C4 (i.e. six people, 4 spots, and order does not matter). This gives you a total of 15 possibilities.

Now, think of the number of ways in which the number of girls will NOT be equal to the number of guys. Two scenarios will results in this.

Scenario 1: 3 girls are selected and one boy is selected

The number of selections that would fit this scenario = 3

because one of three boys can be selected as the "one boy with the three girls".

Scenario 2: 3 boys are selected and one girl is selected

The number of selections that would fit this scenario = 3

because one of three girls can be selected as the "one girl with the three boys".

Hence, the total number of selections in which the number of boys is not equal to the number of girls is 6.

Hence, the number of selections in which the number of boys IS equal to the nunber of girls = 15 (i.e. total number of selection possibilities) - 6

= 9

Hence, the probability of this equal situation = 9/15

I hope this makes sense.
Join the discussion

by somail » Mon Jul 07, 2008 12:22 pm
Thanks, that made a lot of sense. My biggest problem was that I could not figure out how you guys were getting 15. After you mentioned the "formula of combinations", I googled it and found the equation. The problem was much easier after that.
Join the discussion

by ccissilly » Sun Nov 10, 2013 3:33 pm
ok, isnt 6C4 equal to 30? 6*5*4*3*2*1/4*3*2*1?
Join the discussion

by [email protected] » Sun Nov 10, 2013 10:24 pm
Hi ccissilly,

The Combination Formula is:

N! / [K!(N-K)!]

So 6c4 = 6! / [4!(6-4)!] = 6!/4!2! = 15

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by theCodeToGMAT » Sun Nov 10, 2013 11:13 pm
3B 3G --> 4 to be selected

(15 - 3B3*3G1 - 3G3*3B1) = 15 - 1*3 - 1*3 = 9

6c4 = 6x5/2 = 15

Probability = 9/15 = 3/5
R A H U L
Join the discussion

by pareekbharat86 » Mon Nov 11, 2013 9:00 am
Out of 3 girls, choose 2: 3C2= 3

Out of 3 boys, choose 2: 3C2= 3

Therefore, total ways= 3*3 =9

Total ways of choosing 4 persons out of 6= 6C4= 15

Therefore, total probability= 9/15= 3/5
Thanks,
Bharat.
Join the discussion

by gmatclubmember » Mon Nov 11, 2013 9:34 am
Another brute force way is:
We can have a group of 4 either 1B3G, or 2B2G, or 3B1G.
Case1: 3B1G = B1B2B3G1 or B1B2B3G2 OR B1B2B3G3 = total 3 ways
Case3: 3G1B = similar to above = total 3 ways
Case2: 2B2G = {B1B2G1G2, B1B2G2G3, B1B2G1G3} now replace - B1B2 with B2B3 and afterwards with B3B1 (with same G* combinations).
Total combinations will be 3+3+3 = 9

Case 2 is favourable combination: probability = 9/(3+3+9)=9/15=3/5.
As mentioned this is brute force and cannot come in handy with complex combinations.
a lil' Thank note goes a long way :)!!
Join the discussion

by stephny » Tue Nov 12, 2013 12:58 am
please break down your calculation of 6C4 and how that equals 15
Join the discussion

by stephny » Tue Nov 12, 2013 12:59 am
please break down your calculation of 6C4 and how that equals 15
Join the discussion

by theCodeToGMAT » Tue Nov 12, 2013 1:02 am
stephny wrote:please break down your calculation of 6C4 and how that equals 15
Combination formula : nCr ==> n!/r!(n-r)!

So, 6C4 ==> 6!/4!(6-4)! ==> 6!/4!2! ==> 6x5x4!/4!/2! ==> 6x5/2x1 ==> 3x5 ==> 15

I hope it's clear now!
R A H U L
Join the discussion